OpenStudy (anonymous):

how u do this one -10y-6=-2+7y

OpenStudy (phoenixfire):

Add 10y to both sides. Add 2 to both sides. What have you got now?

OpenStudy (anonymous):

17y

OpenStudy (phoenixfire):

and the rest?

OpenStudy (anonymous):

combine like terms.. and solve.

OpenStudy (anonymous):

-10-6=-2+7y +10y +10y --------------- -6=-2+17y +2 +2 --------- 4=17y 4=y

OpenStudy (phoenixfire):

Close but it's -4=17y then divide both sides by 17 \[-{4\over 17}=y\]

OpenStudy (anonymous):

kk thax :)